n-4(1+5)=-2
n-4-20n=-2
4+21n=-2(+4)
21n=2
n=10.5

Answer by checkley71(8403)

N-4(1+5)=-2 THE FIRST OPERATION RERQUIRED IS TO COMBINE THE VALUES INSIDE THE ()
N-4(6)=-2 NOW MULTIPLY THE TERMS
N-24=-2 NOW ADD 24 TO EACH SIDE
N-24+24=-2+24 NOW COMBINE LIKE TERMS.
N=22 ANSWER.
PROOF
22-4(1+5)=-2
22-4(6)=-2
22-24=-2
-2=-2